OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + 2y – z = 7 4x – y + 3z = –2 2x + 2y – z = 9

OpenStudy (anonymous):

Well, here we go: First let's set the first equation equal to: – z = 7 - x - 2y Which becomes: z = x+2y - 7 okay?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

No put this in equation #2: 4x – y + 3z = –2 Becomes: 4x – y + 3(x+2y - 7) = –2 Which becomes: 4x - y +3x +2y -21 = -2

OpenStudy (anonymous):

With me so far?

OpenStudy (anonymous):

Yes kinda

OpenStudy (anonymous):

So, then we combine like terms to get: 4x - y +3x +2y -21 = -2 7x + y -21 = -2 y = -7x +19 And finally put this in the third equation and solve for x. So put in the first and second equations into the third one and solve for x.

OpenStudy (anonymous):

Im not sure

OpenStudy (anonymous):

So we have z = x+2y - 7 and we have y = -7x +19 SO in the third equation: 2x + 2y – z = 9 Wherver you see a z, put that equation, and where you see a y, put in the equation for y. (First put in the y in the z equation)

OpenStudy (anonymous):

I got it! thanks!!

OpenStudy (anonymous):

No problem kiddo.